OpenStudy (anonymous):

Find the equation of the line tangent to y = –x2 + 3x + 8 at x = 2.

OpenStudy (anonymous):

So would it be -x+12?

OpenStudy (er.mohd.amir):

no at x=2 y=10 put at x1=2 and y1=10 and solve

OpenStudy (er.mohd.amir):

x+y=10 is ans.

OpenStudy (anonymous):

It has to be in slope intercept form.

OpenStudy (er.mohd.amir):

y=-x+12 is slope intercept form.

OpenStudy (er.mohd.amir):

diff. wit respect to x dy/dx=-2x+3 at x=2 dy/dx=-1 tagent (y-y1)=(-1)(x-x1)

OpenStudy (anonymous):

Okay thank you!

OpenStudy (er.mohd.amir):

your welcome.